OpenStudy (anonymous):

Solve for y. 5x - y = 3 y = 5x + 3 y = 5x - 3 y = -5x + 3 y = -5x - 3

OpenStudy (anonymous):

Add "y" to both sides. Then subtract 3 from both sides.

OpenStudy (anonymous):

I got D

OpenStudy (anonymous):

If you go step-by-step, using my 2-step method mentioned in my first post, you'll do okay.

OpenStudy (anonymous):

(5x - y) + y = 3 + y 5x + (-y + y) = 3 + y 5x + (0) = 3 + y 5x = 3 + y 5x - 3 = 3 + y - 3 5x - 3 = (3 - 3) + y 5x - 3 = 0 + y 5x - 3 = y

OpenStudy (anonymous):

All good now, @InsanelyChaotic ?

OpenStudy (anonymous):

Are you still stuck, @InsanelyChaotic ? Just look at the last of those 4 lines on the right.

OpenStudy (anonymous):

i can see now:) i had to write it down to understand. thank you